Find the 43rd term of the sequence: 12, 14, 16, 18,... Show all your work.
Answer by prateekagrawal(56)

here first term is 12
and common difference is 14-12= 2
so,
Tn = a+(n-1)d
n= 43, a= 12 and d=2
so, Tn= 12+(43-1)*2 = 12+42*2
Tn= 12+ 84 = 96 is the answer.
so, 43rd term is 96.